Two angles are supplementary. One is 44 degrees more than three time the other. Find the measures of the angles. Supplementary angles are angles whose sum is 180 degrees.
What is the measure of smaller angle? simplify
what is the measure of other angle? simplify

Answer by richwmiller(17219)

a+b=180
a=b+44
2b+44=180
b+22=90
b=68
a=112